# Lampr: Boosting the Effectiveness of Language-Generic Program Reduction via Large Language Models

### 20 Dec 2023 v1

@@ Abstract.
Program reduction is a prevalent technique to facilitate compilers’ debugging by automatically minimizing bug-triggering programs. Existing program reduction techniques are either generic across languages (such as Perses and Vulcan) or specifically customized for one certain language by employing language-specific features, like C-Reduce. However, striking the balance between generality across multiple programming languages and specificity to individual languages in program reduction is an area that still requires exploration.
This paper proposes Lampr, the first LLMs-aided technique utilizing LLMs to perform language-specific program reduction for multiple languages. The core insight is to utilize both the language-generic syntax level program reduction (e.g., Perses and Vulcan) and the language-specific semantic level program transformations learned by LLMs. Alternately, language-generic program reducers efficiently reduce programs into 1-tree-minimality, small enough that is manageable for LLMs; LLMs effectively transform programs via the learned semantics to expose new reduction opportunities for the language-generic program reducers to further reduce the programs.
Our extensive evaluation, encompassing 50 benchmarks across three programming languages (C, Rust, and JavaScript), has highlighted Lampr’s practicality and superiority over Vulcan, the state-of-the-art language-generic program reducer. For effectiveness, Lampr surpasses Vulcan by producing 24.93%, 4.47%, and 11.71% smaller programs on benchmarks in C, Rust and JavaScript, separately. Moreover, Lampr and Vulcan have demonstrated their potential to complement each other. For the C language for which C-Reduce is optimized, by applying Vulcan to the output produced by Lampr, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, Lampr is more efficient when reducing large and complex programs, and takes %, %, % less time than Vulcan to finish all the benchmarks in C, Rust and JavaScript, separately.
@@ 1. Introduction
LLMs-Aided Program Reduction (Lampr). We propose Lampr (Large language model aided program reduction) in this paper, which is, to the best of our knowledge, the first approach that integrates LLMs for program reduction task. Lampr synergistically leverages the strengths of both language-generic program reducers and LLMs. Specifically, Lampr alternates between invoking a language-generic reducer (we use Perses in experiments) and the LLM. Initially, Perses efficiently reduces large programs to a size manageable for the LLM. Subsequently, the LLM further transforms Perses’s output based on specific user-defined prompts that dictate the required transformations. Following this, Perses is re-invoked, as transformations made by the LLM often create additional opportunities for reduction. This process iterates until the program cannot be further minimized. For transformations, we have identified five language-generic transformations to enable further reduction: Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification, and Variable Elimination.
To address the aforementioned challenge of using LLMs, Lampr is designed with a multi-level prompting approach. In detail, Lampr initially requests the LLM to identify a list of potential targets for a given transformation, and then sequentially instructs the LLM to apply the transformation on each target. The multi-level prompt guides the LLM in a more concentrated way, by excluding irrelevant context and other targets that may distract the LLM.
We have conducted extensive evaluations on Lampr, illustrating its superiority over Vulcan, the state-of-the-art language-generic algorithm. On three benchmark suites, i.e., Benchmark-C, Benchmark-Rust and Benchmark-JS, Lampr produces significantly smaller programs than Vulcan by 24.93%, 4.47% and 11.71%, separately. Moreover, Lampr and Vulcan complement each other to some extent. For the C language which C-Reduce is optimized for, by applying Vulcan to the output produced by Lampr, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, Lampr performs comparably to Vulcan. In terms of execution time, Lampr is more efficient than Vulcan on reducing complex programs. Furthermore, our detailed analysis indicates that each of the proposed transformations plays a crucial role in the reduction process. We also compare the performance with the multi-level prompt against that without it, illustrating the efficacy of our proposed multi-level prompting approach.
@@
We introduce LLMs-Aided Program Reduction (Lampr), marking the first algorithm that integrates LLMs into the program reduction process. By synergizing the capabilities of both language-generic tools and LLMs, Lampr achieves a balance between generality across various languages and awareness of semantics in specific languages. Simultaneously, it maintains flexibility in the design and extension of new transformations.
@@
We comprehensively evaluated Lampr on 50 benchmarks across three commonly used languages: C, Rust and JavaScript. Results demonstrate Lampr’s strong effectiveness and generality.
@@ 3. Approach
In this section, we first introduce a motivating example, then we provide an overview of the Lampr workflow. We also outline the details of prompts and proposed transformations in the workflow that enable the LLM to function effectively with given programs.
@@ 3.1. Motivation
A motivating example is displayed in Figure 1, the original code contains highly nested loops, shown in Figure 0(a). From Figure 0(b) to Figure 0(c), the nested loops are fully unrolled into hundreds of lines via Loop Unrolling, based on the semantic transformations from the LLM. Despite the temporary size increase, the following Perses effectively eliminates all lines except for the bug-relevant one. This is also the final result of Lampr, presented in Figure 0(d). By contrast, as shown in Figure 0(e), Vulcan is incapable of escaping the local minima by exhaustively replacing identifiers and tree nodes. Moreover, C-Reduce is not integrated with transformations to unroll loops, and thus cannot fully break down the loop structures. Even though loop unrolling techniques can be added into C-Reduce in future versions, it will be labor-intensive to implement a specific transformation compared to user-defined prompts in natural language.
@@ 3.2. Workflow
The overview of the workflow is outlined in Figure 2. Given a bug-triggering program as input, Lampr invokes a language-generic reducer and the LLM alternately, until the target program cannot be further reduced. In each iteration, the language-generic reducer efficiently reduces the given program to 1-tree-minimality in syntax level. By contrast, the LLM leverages the semantic knowledge of the language and transforms the program given by the language-generic reducer. This process is guided by user-defined prompts, aiming to expose more reduction potentials to the language-generic reducer.
@@
Algorithm 1 shows Lampr’s reduction algorithm. Given as inputs (1) a program targeted for reduction, (2) a property that must be preserved, and (3) the pre-defined prompts, Lampr generates a reduced program . prompts contains primaryQuestion and followupQuestion. primaryQuestion instructs the LLM to identify a list of targets to be transformed, and followupQuestion guides the LLM to apply transformation on each target individually. They will be further introduced in section 3.4. Initially, Lampr loads a sequence of transformations as delineated in Algorithm 1. It then iterates through each transformation, as detailed from Algorithm 1 to Algorithm 1. section 3.3 displays the details of each transformation.
@@
During this process, for each transformation, the algorithm retrieves a predefined primary question along with a follow-up question on Algorithm 1 – Algorithm 1. Lampr first asks the LLM the primary question under the current program. This query aims to guide the LLM to generate a list of specific targets upon which the transformation will be executed. For instance, for Loop Unrolling in the motivating example Figure 0(b), the LLM is asked to identify a list of loops in the given program to be unrolled in the primaryQuestion, and returns a target list .
@@
On Algorithm 1 to 1, Lampr uses followupQuestion to guide LLM to apply the transformation on each identified target within the program. In the motivating example, the followupQuestion can be framed as “Given the program { PROGRAM } and the loop for(i=0;...), optimize it via loop unrolling”. The modified program is then extracted from the LLM’s response text on Algorithm 1. In the example, all loops are unrolled into repeated lines of code in Figure 0(c).
@@
In experimental scenarios, given an input program and the prompt, the LLM may generate multiple transformed programs, as the number of responses can be customized. Among all transformed programs, Lampr keeps the smallest one that still passes the property test, and discards others. If no transformed program returned from this query satisfies the property, Lampr keeps the original one before this query. After each transformation is completed, language-agnostic reducers such as Perses (Sun et al., 2018) are employed to seek additional reduction opportunities, considering that the transformation might have introduced new potentials for further simplification. The algorithm persists in the outermost loop until it reaches a fixpoint, signifying that the program size can no longer be reduced.
@@ 3.3. Transformations
and
denote system prompt
and user prompt provided by the users.
denotes the responses
from the LLM. 3.4. Multi-level Prompts
@@ 4. Evaluation
In this section, we evaluate the effectiveness and efficiency of Lampr. Specifically, we conducted the following research questions.
RQ1. What is the effectiveness of Lampr in program reduction?
RQ2. What is the efficiency of Lampr in program reduction?
RQ3. What is the effectiveness of each transformation in Lampr?
@@ 4.1. Experimental Setup
Within the workflow of Lampr, we employ Perses (Sun et al., 2018) as the language-agnostic reducer due to its superior efficiency compared to Vulcan. Additionally, we utilize OpenAI API (ope, 2023a), specifically the gpt-3.5-turbo-0613 version, to serve as the LLM. We also develop a variant named LPR+Vulcan, which invokes Vulcan to further reduce the program after Lampr finishes. For a fair comparison, all algorithms were executed in a single-process, single-thread environment.
Benchmarks. To measure the effectiveness and efficiency of Lampr across various languages, we employ three benchmark suites: Benchmark-C, Benchmark-Rust and Benchmark-JS. The Benchmark-C, previously collected and utilized by previous studies (Xu et al., 2023; Tian et al., 2023c; Zhang et al., 2023), comprises 20 large complex programs triggering real-world bugs in LLVM or GCC. Benchmark-Rust, incorporating 20 bug-triggering Rust programs, has also been used in prior research (Xu et al., 2023). We further craft Benchmark-JS, a non-public benchmark suite, for this study. Specifically, we use FuzzJIT (Wang et al., 2023) to fuzz a prevalent JavaScript engine, i.e., JavaScriptCore (version c6a5bcc), and then randomly collect 10 programs that cause miscompilations in JIT compiler. Since the programs and reduced programs in Benchmark-JS are not publicly available and thus not in the training sets of LLMs. The evaluation on Benchmark-JS helps us investigate whether Lampr suffers from the data leakage problem (Wu et al., 2023). In total, the evaluation benchmarks encompass 50 programs triggering real-world bugs in compilers, spanning across three popular programming languages.
@@ 4.2. RQ1: Effectiveness
We measure the effectiveness of Lampr, LPR+Vulcan and baseline algorithms via the final program size in token. A smaller size is favored, as it signifies the removal of more bug-irrelevant code, thereby saving developers more manual effort. The effectiveness of each algorithm on all three benchmark suites is presented in Table 1. Due to the randomness of Lampr, we repeat five times for Lampr and LPR+Vulcan on every benchmark, and display the mean and standard deviation in the table. On each benchmark, the minimal results are highlighted in bold.
| Benchmark | Original | Perses | Vulcan | C-Reduce | Lampr | Lampr+Vulcan | |||
@@
Benchmark-C. On this benchmark suite, Perses reduces the programs to an average of 247.8 tokens. Building upon this, Vulcan further compresses the average program size into 157.4, i.e., thereby continues to decrease the program size by 34.35%. Despite Vulcan’s notable reduction progress, Lampr is still capable of continuing to push the limit of Perses, and reduces the programs in Benchmark-C into 105.8 tokens on average across all five runs. It cuts down the average program size of Perses by 51.33%, outperforming Vulcan significantly by 24.93% (proved by a p-value of 0.002). C-Reduce stands out by achieving the lowest average program size, i.e., 85.7 tokens. This performance is anticipated as C-Reduce incorporates various transformation passes specifically designed for C. Despite relying on only general transformations, LPR+Vulcan still achieves performance comparable to C-Reduce, averaging 86.1 tokens. Moreover, it outperforms C-Reduce in 13 out of 20 benchmarks, highlighting its effectiveness with only language-generic transformations.
@@
Benchmark-Rust. On Benchmark-Rust, Perses and Vulcan produce programs with 212.5 and 184.2 tokens on average, separately. Lampr and LPR+Vulcan further shrink the average program size into 147.5 and 135.5 tokens. C-Reduce produces the second-largest programs on average, only smaller than Perses. This is anticipated since C-Reduce lacks specialized transformations for Rust.
@@
Further analysis into these benchmarks reveals that Vulcan and transformations in Lampr are complementary on Benchmark-Rust. Vulcan shows higher effectiveness on reducing relatively smaller programs, as evidenced by the average original size of 43 tokens in the 9 programs where it is better than Lampr. In contrast, the 15 programs where Lampr is proved better than Vulcan have an average of 269 tokens, indicating its proficiency on reducing relatively larger programs. Our speculation is that Vulcan and Lampr target different reduction opportunities. Vulcan performs identifier/sub-tree replacement and local exhaustive search. Such reducers, while lacking in semantic analysis, find reduction opportunities in a ”brute-force” manner and is particularly effective at uncovering less obvious reduction opportunities. On the other hand, Lampr employs more semantic and intelligent transformations, adeptly and systematically analyzing and reducing a complex program step by step. Moreover, results of LPR+Vulcan in the last column prove the complementary characteristic between Lampr and Vulcan, which achieve the best in 16 out of the total 20 benchmarks.
@@
Benchmark-JS. Programs in Benchmark-JS are much simpler and smaller than those in the previous two benchmark suites. Therefore, even Perses alone is capable of reducing the programs to only 55.5 tokens. Following this, Vulcan, Lampr and LPR+Vulcan achieve 38.2, 33.9 and 27.5 tokens, further reducing the average results by 30.35%, 38.66% and 38.66%, separately. Similar to its performance on Rust, C-Reduce cannot outperform the aforementioned algorithms on JavaScript, due to its lack of employment of JavaScript’s semantics. The evaluation results also serve to demonstrate that the performance exhibited by Lampr is not attributable to data leakage. These benchmarks were collected by the authors via fuzzing, and the optimal results remain inaccessible to the public, thereby precluding any possibility of LLMs memorizing them.

### 02 Mar 2024 v2

# LPR: Large Language Models-Aided Program Reduction

@@ Abstract.
Program reduction is a widely used technique to facilitate debugging compilers by automatically minimizing programs that trigger compiler bugs. Existing program reduction techniques are either generic to a wide range of languages (such as Perses and Vulcan) or specifically optimized for one certain language by exploiting language-specific knowledge (e.g., C-Reduce). However, synergistically combining both generality across languages and optimality to a specific language in program reduction is yet to be explored.
This paper proposes LPR, the first LLMs-aided technique leveraging LLMs to perform language-specific program reduction for multiple languages. The key insight is to utilize both the language generality of program reducers such as Perses and the language-specific semantics learned by LLMs. Concretely, language-generic program reducers can efficiently reduce programs into a small size that is suitable for LLMs to process; LLMs can effectively transform programs via the learned semantics to create new reduction opportunities for the language-generic program reducers to further reduce the programs.
Our thorough evaluation on 50 benchmarks across three programming languages (i.e., C, Rust and JavaScript) has demonstrated LPR’s practicality and superiority over Vulcan, the state-of-the-art language-generic program reducer. For effectiveness, LPR surpasses Vulcan by producing 24.93%, 4.47%, and 11.71% smaller programs on benchmarks in C, Rust and JavaScript, separately. Moreover, LPR and Vulcan have the potential to complement each other. For the C language for which C-Reduce is optimized, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR is more efficient when reducing large and complex programs, taking %, %, % less time than Vulcan to finish all the benchmarks in C, Rust and JavaScript, separately.
@@ 1. Introduction
LLMs-Aided Program Reduction (LPR). We propose LPR (Large language model aided program reduction) in this paper, which is, to the best of our knowledge, the first approach that integrates LLMs for program reduction task. LPR synergistically leverages the strengths of both language-generic program reducers and LLMs. Specifically, LPR alternates between invoking a language-generic reducer (we use Perses in experiments) and the LLM. Initially, Perses efficiently reduces large programs to a size manageable for the LLM. Subsequently, the LLM further transforms Perses’s output based on specific user-defined prompts that dictate the required transformations. Following this, Perses is re-invoked, as transformations made by the LLM often create additional opportunities for reduction. This process iterates until the program cannot be further minimized. For transformations, we have identified five language-generic transformations to enable further reduction: Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification, and Variable Elimination.
To address the aforementioned challenge of using LLMs, LPR is designed with a multi-level prompting approach. In detail, LPR initially requests the LLM to identify a list of potential targets for a given transformation, and then sequentially instructs the LLM to apply the transformation on each target. The multi-level prompt guides the LLM in a more concentrated way, by excluding irrelevant context and other targets that may distract the LLM.
We have conducted extensive evaluations on LPR, illustrating its superiority over Vulcan, the state-of-the-art language-generic algorithm. On three benchmark suites, i.e., Benchmark-C, Benchmark-Rust and Benchmark-JS, LPR produces significantly smaller programs than Vulcan by 24.93%, 4.47% and 11.71%, separately. Moreover, LPR and Vulcan complement each other to some extent. For the C language which C-Reduce is optimized for, by applying Vulcan to the output produced by LPR, we can attain program sizes that are on par with those achieved by C-Reduce. For efficiency, LPR performs comparably to Vulcan. In terms of execution time, LPR is more efficient than Vulcan on reducing complex programs. Furthermore, our detailed analysis indicates that each of the proposed transformations plays a crucial role in the reduction process. We also compare the performance with the multi-level prompt against that without it, illustrating the efficacy of our proposed multi-level prompting approach.
@@
We introduce LLMs-Aided Program Reduction (LPR), marking the first algorithm that integrates LLMs into the program reduction process. By synergizing the capabilities of both language-generic tools and LLMs, LPR achieves a balance between generality across various languages and awareness of semantics in specific languages. Simultaneously, it maintains flexibility in the design and extension of new transformations.
@@
We comprehensively evaluated LPR on 50 benchmarks across three commonly used languages: C, Rust and JavaScript. Results demonstrate LPR’s strong effectiveness and generality.
@@ 3. Approach
In this section, we first introduce a motivating example, then we provide an overview of the LPR workflow. We also outline the details of prompts and proposed transformations in the workflow that enable the LLM to function effectively with given programs.
@@ 3.1. Motivation
A motivating example is displayed in Figure 1, the original code contains highly nested loops, shown in Figure 0(a). From Figure 0(b) to Figure 0(c), the nested loops are fully unrolled into hundreds of lines via Loop Unrolling, based on the semantic transformations from the LLM. Despite the temporary size increase, the following Perses effectively eliminates all lines except for the bug-relevant one. This is also the final result of LPR, presented in Figure 0(d). By contrast, as shown in Figure 0(e), Vulcan is incapable of escaping the local minima by exhaustively replacing identifiers and tree nodes. Moreover, C-Reduce is not integrated with transformations to unroll loops, and thus cannot fully break down the loop structures. Even though loop unrolling techniques can be added into C-Reduce in future versions, it will be labor-intensive to implement a specific transformation compared to user-defined prompts in natural language.
@@ 3.2. Workflow
The overview of the workflow is outlined in Figure 2. Given a bug-triggering program as input, LPR invokes a language-generic reducer and the LLM alternately, until the target program cannot be further reduced. In each iteration, the language-generic reducer efficiently reduces the given program to 1-tree-minimality in syntax level. By contrast, the LLM leverages the semantic knowledge of the language and transforms the program given by the language-generic reducer. This process is guided by user-defined prompts, aiming to expose more reduction potentials to the language-generic reducer.
@@
Algorithm 1 shows LPR’s reduction algorithm. Given as inputs (1) a program targeted for reduction, (2) a property that must be preserved, and (3) the pre-defined prompts, LPR generates a reduced program . prompts contains primaryQuestion and followupQuestion. primaryQuestion instructs the LLM to identify a list of targets to be transformed, and followupQuestion guides the LLM to apply transformation on each target individually. They will be further introduced in section 3.4. Initially, LPR loads a sequence of transformations as delineated in Algorithm 1. It then iterates through each transformation, as detailed from Algorithm 1 to Algorithm 1. section 3.3 displays the details of each transformation.
@@
During this process, for each transformation, the algorithm retrieves a predefined primary question along with a follow-up question on Algorithm 1 – Algorithm 1. LPR first asks the LLM the primary question under the current program. This query aims to guide the LLM to generate a list of specific targets upon which the transformation will be executed. For instance, for Loop Unrolling in the motivating example Figure 0(b), the LLM is asked to identify a list of loops in the given program to be unrolled in the primaryQuestion, and returns a target list .
@@
On Algorithm 1 to 1, LPR uses followupQuestion to guide LLM to apply the transformation on each identified target within the program. In the motivating example, the followupQuestion can be framed as “Given the program { PROGRAM } and the loop for(i=0;...), optimize it via loop unrolling”. The modified program is then extracted from the LLM’s response text on Algorithm 1. In the example, all loops are unrolled into repeated lines of code in Figure 0(c).
@@
In experimental scenarios, given an input program and the prompt, the LLM may generate multiple transformed programs, as the number of responses can be customized. Among all transformed programs, LPR keeps the smallest one that still passes the property test, and discards others. If no transformed program returned from this query satisfies the property, LPR keeps the original one before this query. After each transformation is completed, language-agnostic reducers such as Perses (Sun et al., 2018) are employed to seek additional reduction opportunities, considering that the transformation might have introduced new potentials for further simplification. The algorithm persists in the outermost loop until it reaches a fixpoint, signifying that the program size can no longer be reduced.
@@ 3.3. Transformations
3.4. Multi-level Prompts
@@ 4. Evaluation
In this section, we evaluate the effectiveness and efficiency of LPR. Specifically, we conducted the following research questions.
RQ1. What is the effectiveness of LPR in program reduction?
RQ2. What is the efficiency of LPR in program reduction?
RQ3. What is the effectiveness of each transformation in LPR?
@@ 4.1. Experimental Setup
Within the workflow of LPR, we employ Perses (Sun et al., 2018) as the language-agnostic reducer due to its superior efficiency compared to Vulcan. Additionally, we utilize OpenAI API (ope, 2023a), specifically the gpt-3.5-turbo-0613 version, to serve as the LLM. We also develop a variant named LPR+Vulcan, which invokes Vulcan to further reduce the program after LPR finishes. For a fair comparison, all algorithms were executed in a single-process, single-thread environment.
Benchmarks. To measure the effectiveness and efficiency of LPR across various languages, we employ three benchmark suites: Benchmark-C, Benchmark-Rust and Benchmark-JS. The Benchmark-C, previously collected and utilized by previous studies (Xu et al., 2023; Tian et al., 2023c; Zhang et al., 2023), comprises 20 large complex programs triggering real-world bugs in LLVM or GCC. Benchmark-Rust, incorporating 20 bug-triggering Rust programs, has also been used in prior research (Xu et al., 2023). We further craft Benchmark-JS, a non-public benchmark suite, for this study. Specifically, we use FuzzJIT (Wang et al., 2023) to fuzz a prevalent JavaScript engine, i.e., JavaScriptCore (version c6a5bcc), and then randomly collect 10 programs that cause miscompilations in JIT compiler. Since the programs and reduced programs in Benchmark-JS are not publicly available and thus not in the training sets of LLMs. The evaluation on Benchmark-JS helps us investigate whether LPR suffers from the data leakage problem (Wu et al., 2023). In total, the evaluation benchmarks encompass 50 programs triggering real-world bugs in compilers, spanning across three popular programming languages.
@@ 4.2. RQ1: Effectiveness
We measure the effectiveness of LPR, LPR+Vulcan and baseline algorithms via the final program size in token. A smaller size is favored, as it signifies the removal of more bug-irrelevant code, thereby saving developers more manual effort. The effectiveness of each algorithm on all three benchmark suites is presented in Table 1. Due to the randomness of LPR, we repeat five times for LPR and LPR+Vulcan on every benchmark, and display the mean and standard deviation in the table. On each benchmark, the minimal results are highlighted in bold.
Benchmark | Original | Perses | Vulcan | C-Reduce | LPR | LPR+Vulcan | |||
@@
Benchmark-C. On this benchmark suite, Perses reduces the programs to an average of 247.8 tokens. Building upon this, Vulcan further compresses the average program size into 157.4, i.e., thereby continues to decrease the program size by 34.35%. Despite Vulcan’s notable reduction progress, LPR is still capable of continuing to push the limit of Perses, and reduces the programs in Benchmark-C into 105.8 tokens on average across all five runs. It cuts down the average program size of Perses by 51.33%, outperforming Vulcan significantly by 24.93% (proved by a p-value of 0.002). C-Reduce stands out by achieving the lowest average program size, i.e., 85.7 tokens. This performance is anticipated as C-Reduce incorporates various transformation passes specifically designed for C. Despite relying on only general transformations, LPR+Vulcan still achieves performance comparable to C-Reduce, averaging 86.1 tokens. Moreover, it outperforms C-Reduce in 13 out of 20 benchmarks, highlighting its effectiveness with only language-generic transformations.
@@
Benchmark-Rust. On Benchmark-Rust, Perses and Vulcan produce programs with 212.5 and 184.2 tokens on average, separately. LPR and LPR+Vulcan further shrink the average program size into 147.5 and 135.5 tokens. C-Reduce produces the second-largest programs on average, only smaller than Perses. This is anticipated since C-Reduce lacks specialized transformations for Rust.
@@
Further analysis into these benchmarks reveals that Vulcan and transformations in LPR are complementary on Benchmark-Rust. Vulcan shows higher effectiveness on reducing relatively smaller programs, as evidenced by the average original size of 43 tokens in the 9 programs where it is better than LPR. In contrast, the 15 programs where LPR is proved better than Vulcan have an average of 269 tokens, indicating its proficiency on reducing relatively larger programs. Our speculation is that Vulcan and LPR target different reduction opportunities. Vulcan performs identifier/sub-tree replacement and local exhaustive search. Such reducers, while lacking in semantic analysis, find reduction opportunities in a ”brute-force” manner and is particularly effective at uncovering less obvious reduction opportunities. On the other hand, LPR employs more semantic and intelligent transformations, adeptly and systematically analyzing and reducing a complex program step by step. Moreover, results of LPR+Vulcan in the last column prove the complementary characteristic between LPR and Vulcan, which achieve the best in 16 out of the total 20 benchmarks.
@@
Benchmark-JS. Programs in Benchmark-JS are much simpler and smaller than those in the previous two benchmark suites. Therefore, even Perses alone is capable of reducing the programs to only 55.5 tokens. Following this, Vulcan, LPR and LPR+Vulcan achieve 38.2, 33.9 and 27.5 tokens, further reducing the average results by 30.35%, 38.66% and 38.66%, separately. Similar to its performance on Rust, C-Reduce cannot outperform the aforementioned algorithms on JavaScript, due to its lack of employment of JavaScript’s semantics. The evaluation results also serve to demonstrate that the performance exhibited by LPR is not attributable to data leakage. These benchmarks were collected by the authors via fuzzing, and the optimal results remain inaccessible to the public, thereby precluding any possibility of LLMs memorizing them.